A quality control engineer is interested in engineering the proportion of defective items coming off a production line, in a sample of 300 items. in 27 are defective. Calculate a 90% confidence interval for the proportion of defective from this production.
Answer by ewatrrr(24785)

p = 27/300 = .09, mean = 27, SD = sqrt(27*.91) = 4.957
ME = 1.645*4.957/sqrt(300) = .47 0r .5
CI: μ-ME < u < μ + ME
CI: (26.5,27.5)